OpenStudy (anonymous):

Solve. 4x + 3 = 47 x = 11* x = 40 x = 44 x = 50 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

4x + 3 = 47 => 4x = 47 - 3 = 44 => x = 11

OpenStudy (anonymous):

4x + 3 = 47 Subtract 3 from both sides: 4x = 43 Divide: x = 10.75

OpenStudy (anonymous):

@AashishKumar6 47 - 4 = 43 not 44.

OpenStudy (anonymous):

But the answer is still 11 because you have to round 10.75 to 11.

OpenStudy (anonymous):

@greenlegodude57 It's 47 - 3 = 44.

OpenStudy (anonymous):

Ooops, my bad.